Please help
A car travels from one town to another at a speed of 42mph. If it had gone 6 mph faster, it could have made the trip in a half hour less time. How far apart are the towns?

Answer by rwm(914)

42x=48(x-.5)
42x=48x-24
6x=24
x=4
42*4=168 miles apart
check
42*4=48*(3.5)
168=168